OpenStudy (anonymous):

Hey I have a question with EQUATIONS BY COMPLETING THE SQUARES: x^2--10x+6=0 my answer was x=+or- Root 19 + 5 therefore X=5root19 or 4root 19? is that right? Hey I have a question with EQUATIONS BY COMPLETING THE SQUARES: x^2--10x+6=0 my answer was x=+or- Root 19 + 5 therefore X=5root19 or 4root 19? is that right? @Mathematics

OpenStudy (anonymous):

* i meant and

OpenStudy (anonymous):

x^2 -10x = -6 x^2-10x + 25 = 25 -6 (x-5)^2=19 x= 5 +-root19

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

i got another one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay so i started it off but i got stuck on how to finish

OpenStudy (anonymous):

it is x^2-5x+1=0 so far i got x^2-5x+25/4=-1(25/4)

OpenStudy (anonymous):

what i want to know is how to Write the left side as the square of a binomial.

OpenStudy (anonymous):

@mathwiz707